UT Jul 13 - Extra Engineering Night - rigid installed on SX



Night Info

  • Sunset 19:29
  • Nautical Twilight Ends 20:32
  • Astro Twilight Ends 21:08
  • Astro Twilight Starts 03:43
  • Nautical Twilight Starts 04:19
  • Sunrise 05:22
  • Moon Illumination 99% - rise and set before and after nautical twilights

Staff

  • OSA: D.G.Huerta
  • ISp: Mtn staff
  • ISA: B.Rothberg
  • Mtn Manager:B.Bingham

Issues

  • MODS2B sequencer - all but 2 images affected - IT 6386
  • MODS2B dewar pressure jumped from 1e-6 to 1e-3 after HEB restart - IT 8679
  • Had to "rejiggle" the IMCS lasers on MODS2
  • QTH1 lamp on MODS1 appears to be burnt out - IT 8660


Closed Dome

  • DONE MODS calibrations for LBTB - 2hr
    • ~/Service_Observing/2022A_D_E/MODS/Bosman_MODS_calibrations
  • DONE TMS testing with MODS in the beam - check for impact on MODS observations - 1hr?
  • Closed dome MODS cals
    • DONE Min: 1-2 flats in each mode and each filter (to look for debris on the optics).
  • Direct mode laser on/laser off (lbto@robs:~/kuhn/).
    • You would need to execute the commands "irlaser disable", "irlaser off' in the command window.
    • The scripts do not do this. Run the disabled after the enabled (that way the collimator is aligned and at the same position as for the laser on images).
      • laser_enabled_blue.obs

      • laser_disabled_off_blue.obs

      • laser_enabled_red.obs

      • laser_disabled_off_red.obs


Night Plan

  • Goal:
    • Obtain absolute flux comparison among SX w/rigid and DX w/ASM and SX w/ASM - requires photometric conditions
    • If non-photometric conditions - then relative flux comparisons using spectrophotometric stars and photometric standards
    • Use f/15 instruments (MODS, PEPSI, LUCI)
    • See IT 8600
    • Also see IT 7730
  • Starting Configuration: bino-MODS - highest priority given forecast
    • then PEPSI, then LUCI (in order of priority)
  • Short Verificaiton rigid is installed on SX
    • Successful preset with guiding/WFS'ing on AGWs
    • stars are round
  • MODS Spectrophotometric standards
    • About 1 hr each for grating, prism.
    • Scripts: ~/Service_Observing/2022A_D_E/MODS/2022Jul13/
    • Observe 2 standards - different airmasses use dual, blue, and red gratings - highest priority
      • BD+33 2642 15 51 59 +32 56 54 - twilight-01:45hr - 60,75 deg from Moon
        • BD+33 observed a few days ago in clear conditions - 20220707
      • Feige110 23 19 58 -05 09 56 - visible 01hr-twlight - 78, 60deg from Moon
    • Photometric standards (all sloan filters)
      • M13 16 41 41 +36 27 40 - visible twlight-01:45 - 60-70 deg from Moon
      • M92 17 17 07 +43 08 09 - visible twilight-03:45hr - 73-79 deg from Moon
      • M71 19 53 46 +18 46 45 - visible all night - 53-48 deg from Moon
    • Prism mode for spectrophotometric standards (if clear)
      • Feige110 23 19 58 -05 09 56 - visible 01hr-twlight
        • flux tables for both 10-A and 50-A bins.
  • PEPSI Flux Standards Available for the Night
    • HR 5501 14 45 30 +00 43 02 V=5.6 - visibile twilight-23:15hr
    • HR 7596 19 54 44 +00 16 25 V=5.6 -visible 21:15-twilight
    • HR 7950 20 47 40 -09 29 44 V=3.7 - visible 22:45-twilight
    • HR 153 00 36 58 +53 53 48 V=3.6 - visible midnight-twilight
    • Restriced blue CD setting to avoid hardware issues
  • LUCI Photometric standards
    • use J,H, and K both sides - relative or absolute flux
      • Y1 comparisons - cutoff is 0.95 micron on blue-end - matches LUCI-2 dichroic cutoff
    • OT file: /home/lbto/Service_Observing/2022A_D_E/PoorWeather/LUCI_Phot_Standards/LUCI_Phot_ZP_OT_V5.xml
    • Standards available all night
  • Other (tasks lower priority to try once above are completed)
    • MODS trailed images issue


Night Log

Closed Dome Activities

  • UT 01:42 - MODS GUIs up - if we go to zenith will wake MODS, bringing up LUCI GUIs too. Won't init all until we move to zenith
  • UT 01:52 - all-sky is very grey and swirly
  • UT 02:38 - I see virga on the roof cam
  • UT 02:46 - planning to move back to zenith
  • UT 03:31 - telescope at zenith - reconfiguring for bino-MODS - waking MODS now
  • UT 03:49 - taking maskSnaps - checking new masks inserted
  • UT 03:52 - had to abort MODS2 got double IMCS lock failure (timed out after 120sec, when I hit retry it imeediately popped up retry - so CTRL C'd)
  • AllSkyCurrentImage.JPG
  • UT 03:55 - MODS-2 IMCS lock failed again
    • putting MODS-2 to sleep and waking again
    • typing irlaser in MODS2 GUI - IRPOUTPUT = 0 - time to troubleshoot
    • irlaser disable
    • irlaser off isisCmd --mods2 m2.ie irlaser off and then wait...
    • isisCmd --mods2 m2.ie irlaser on
    • isisCmd --mods2 m2.ie irlaser enable
    • isisCmd --mods2 m2.ie irlaser --> IRPOUT=1.2
  • UT 04:04 - trying maskSnap on MODS-2 again
  • UT 04:10 - maskSnap complete - both look OK on both MODS
  • UT 04:15 - updated IT 6386 - bad sequencer MODS2 Blue - last entry in the IT was 2020! - file showing problem is mods2b.20220713.0002.fits
  • UT 04:39 - sky still looks bleak, but the space alien moved closer to LBT
  • UT 09:18 - all-sky now looks worse - had thinned a bit earlier getting our hopes up, but now, not so much
TMS Tests with MODS
  • Configuring for test:
    • Putting MODS into Observing Mode
    • putting 2.4" longslit into focal plane
    • instconfig dual grating (on GUI commend line)
    • lcoking IMCS manually on each channel
    • manually set to 1800sec per channel
  • UT 04:25 - starting test exposures (modsN.20220713.XXXX)
  • last test saw an emission line on MODS1 only at 830 nm
  • M2B 2nd exposure shows sequencer issue as well (mods2b.20220713.0003.fits)
  • Do see emission lines in both M1R and M2R - (X=3326,3156)
  • Next step, take some exposure length with laser off
  • emission line on M1R still there with TMS laser off, also there on M2R (still fainter)
  • fourth image on M2B sequencer also shows issue - updating IT 6386
Troubleshooting M2B sequencer (and then some)
  • Going to troubleshoot sequencer issue by restarting the M2B IC
    • UT 05:39 - restarted M2B IC - taking another 30 sec exposure to test
    • first image mods2b.20220713.0005.fits looks good
    • M2B 0006 bad, 0007 bad, 0008 bad, - each one is slighlty different sequencing shift
    • M2B 0009, 0010 bad
    • All five biases are bad
    • UT 05:54 - Now rebooting M2B iC with ctrl-alt-delete
    • restarting IC after reboot
    • taking another 5 biases just on MODS-2
    • rebooting and restarting IC did not fix the issue
    • M2B 0011-0015 - only 0013 looks OK
    • Next step - restart the HEB - will do once biases finish
    • UT 06:04 - restarting HEB (M2B)
    • UT 06:10 - M2B vaccum now showing 5.4-e-3 torr after power cycling HEBs - wasn't soft a few minutes ago and was e-6 torr during afternoon instruments update
    • UT 06:12 - now 7.6e-3 torr, temp is staying the same during this
    • O.Kuhn calling E.Solheid - he suspects the pressure sensor may have gone bad since the CCD temp has remained constant
    • UT 06:22 - now at 9e-3 torr but temp remains at -109C - another option is to pump on M2 dewar
    • UT 06:24 - O.Kuhn will call R.Pogge - left message
    • Submitted pressure change as IT 8679
    • UT 06:48 -cycling IG power - no change
    • UT 07:51 - vacuum says 1e-02 torr CCD temp -109.1C
LBTB Bosman Cals for MOS masks
  • filenames: modsN.20220713.XXXX
  • UT 06:50 - red grating only cals - should not be impacted by M2B issues
    • ID 521532
    • ID 551224
      • MODS-1 lamp issue
        • M1R 0025-0029 - QTH1 not firing? blank looks like bias level counts
        • UT 07:30 - re-sending script for MODS1-red
        • M1R 0030 - still no flux
        • IT 8680
      • M1R 0031-0033 - trying QTH2 to see if any flux - worked, altering scripts - these are probably OK to use - but better to use ones where script has header correct
      • M1R 0034-0048
      • M2R 0030-0044

Red Only Imaging Flats
  • UT 08:37 - running scripts for flats (red-only) - r,i,z filters
    • MODS-1 looks like something bottom right - field lens
    • MODS-2 also see something bottom right
    • M1R.pngM2R.png
    • 5th exposure of z-band M1R - stuck on the long readout
    • M1R 0049-0063
    • M2R 0045-0059
  • UT 10:54 - logged out of LUCI and PEPSI computers
  • UT 11:15 - let's stick a fork in it and call it done...

-- BarryRothberg - 04 Jul 2022
I Attachment Action Size Date Who Comment
AllSkyCurrentImage.JPGJPG AllSkyCurrentImage.JPG manage 58 K 13 Jul 2022 - 03:54 BarryRothberg  
M1R.pngpng M1R.png manage 174 K 13 Jul 2022 - 08:43 BarryRothberg  
M2R.pngpng M2R.png manage 189 K 13 Jul 2022 - 08:43 BarryRothberg  
MODS_img_phot_12Jul2022.pngpng MODS_img_phot_12Jul2022.png manage 21 K 08 Jul 2022 - 17:15 OlgaKuhn Visibility chart for MODS photometric imaging fields (12/13 Jul 2022)
MODS_specphot_vis_12Jul2022.pngpng MODS_specphot_vis_12Jul2022.png manage 17 K 08 Jul 2022 - 17:12 OlgaKuhn Visibility chart for MODS spectrophotometric standards (12/13 Jul 2022)
Topic revision: r11 - 13 Jul 2022, BarryRothberg
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback